# Implications of computer science theory for the simulation hypothesis

### v2

@@ 2.1 Notation
My notation is conventional. The set of of all positive integers is ${\mathbb{N}}$, and the set of non-negative integers is +=${\mathbb{Z}}^{+}={\mathbb{N}}\cup\{0\}$. I write $|X|$ for the cardinality of any set $X$. In addition, for any set $X$ I write $X^{*}$ for the set of all finite strings of elements of $X$. Note that so long as $X$ is finite, $X^{*}$ is countable. As an important example, ${\mathbb{B}}^{*}$ is the set of all finite bit strings. Since that set can be bijectively mapped to ${\mathbb{N}}$, I will follow convention and treat finite bit strings as positive integers and vice-versa, with the bijection implicit.
As discussed in Appendix A, I write $T^{m}(x)$ for the (possibly partial) function given by running the TM with index $m$ on input $x$ until it halts (with the function undefined for input $x$ if it does not halt for that input). So $m$ ranges over all natural numbers ${\mathbb{N}}$. Often I assume, implicitly or otherwise, that certain quantities are in the form of prefix-encoded bit strings. In particular, I use some standard bijective encoding function of all tuples of bit strings into a single bit string, indicated using angle brackets, .,.,.,.,$\langle$.,$\rangle$,$\langle$.,.,$\rangle$ . , . ⟩ , ⟨ . , . , . ⟩, etc. I assume that this encoding is non-decreasing in the number of arguments, i.e., for any finite set of bit strings $\{b(1),b(2),\ldots,b(m)\}$, the length of the bit string $\langle b(1),b(2),\ldots,b(m-1)\rangle$ is not greater than that of $\langle b(1),b(2),\ldots,b(m)\rangle$.
Other important definitions and notations is in Appendix A , e.g., of partial functions, and of computable functions (here always assumed to be total). I also review the definition of “computational universality” in that appendix. In addition, there I review the definitions of a universal TM (UTM), and a prefix-free TM (the implementation of TMs I will often be assume in this paper, implicitly or otherwise). I also define the instantaneous description (ID) of a Turing machine (TM) there. In the main body of this paper I will assume that the non-blank alphabet of the TM, $\Lambda$, is just ${\mathbb{B}}$. So as described in Appendix A, we can take the state of the tape to be a finite string in ${\mathbb{B}}^{*}$, even though strictly speaking the definition of TMs assumes infinitely long tapes.
@@ 2.2 Framework for analyzing CS theory of the simulation hypothesis
I will also mostly be interested in cases where the computer simulates the evolution of the state of the universe external to the computer (the environment of the computer) and/or its own evolution. This means that the state space of the environment must be finite (though arbitrarily large), to ensure that its state at a particular time can be appropriately encoded on the input of a UTM.
This leads me to write the state space of a (computational) universe as
@@
where $W$ is initially finite and cannot be extended dynamically, whereas $N$ is countably infinite. I will parameterize $N$ as $\underline{N}\times R$.
In this decomposition, $R$ is a finite set that represents the internal variables of the computer. So for example, in a conventionally represented TM, $R$ would be the state of the TM’s head, the position(s) of its pointer(s) on the tape(s), etc.  [^5] Strictly speaking, in a TM the pointer variables can have any value in ${\mathbb{N}}$, since they can point to an arbitrary position on the (countably infinite) tape(s). This technical concern can be addressed several ways. For example, we might want to have the states of the pointer variables at any iteration after the $k$’th iteration given by the state of some special subsection of the tape(s), $\underline{n}$, with the states of those variables at earlier iterations recorded in $R$. Another possibility would be to have $R$ infinitely extendable.
$\underline{N}={\mathbb{B}}^{*}$ is instead the state of the tape(s) in the case that the computer is represented as a TM, or it is the state of the memory in the case that the computer is represented as a RAM machine, or an appropriately modified laptop computer. The reason for writing the countably infinite space $\underline{N}$ as the set of finite bit strings is so that we can physically implement it as a system that initially has a finite state space, but whose state space can be expanded by an arbitrarily large finite amount an arbitrary number of times as the universe evolves. See Section 2.6. I will write the elements of $\underline{N}$ as $\underline{n}$.
@@
To simplify the analysis, I do not assume that the models of physical universes that I investigate in this paper capture those universes in toto, e.g., if the state spaces of those entire universes are actually uncountably infinite. Similarly, I will not assume that a simulation running in a computer inside a universe models the dynamics of an entire universe. Rather throughout this paper the expression “universe” should be understand as shorthand for “possibly coarse-grained subset of a universe”. However I do assume that the universe obeys deterministic dynamics, be it over the original state space (assuming it is countable) or some coarse-grained state space. As an example, in Section 2.6 I present a detailed example of what such a “coarse-grained subset of a universe” could be for our particular physical universe.
I write the evolution of the universe from an initial, time-00 state $(w_{0},n_{0})$ to a time $t$ state as a vector-valued evolution function $g$ of the initial state of that universe,
@@
Note that the image of $g$ is the Cartesian product $W\times N=W\times(\underline{N}\times R)$. I use usual notation for components of vector-valued functions. So in particular, $g(t,w_{0},n_{0})_{\underline{N}}$ is the $\underline{N}$-component of $g(t,w_{0},n_{0})$.
@@ 2.3 What it means for one universe to simulate another
The term “simulation” was not given a formal definition in any of the previous literature on the simulation hypothesis. Moreover, “simulation” (and the associated term “bisimulation”) already has a formal definition in the CS theory of state transition systems [63]. However, despite its name, this definition from CS theory does not describe what “simulation” is loosely understood to mean in the context of the simulation hypothesis.
In this paper $V$ “simulates” $V^{\prime}$ if for all initial states $v^{\prime}_{0}$ of $V$, and all amounts of time $\Delta t^{\prime}$ into the future of $V^{\prime}$ that we might to simulate the dynamics of $V^{\prime}$, there is some associated initial state of $N$ and of $W$, together with a time into the future of the simulating universe, $\Delta t$, such that the state of the computer $N$ at that future time $\Delta t$ gives us the desired state of $V^{\prime}$ at its future time $\Delta t^{\prime}$.
I formalize this as follows:
@@
Note that the LHS of Eq. 3 is the second of the two components of the vector-valued function $g$, while the RHS is an encoding of both components of $g^{\prime}$ into a single variable. Note also the requirement in Definition 1 that Eq. 3 hold for all $t\geq\tau$, and so the state of the simulating computer $N$ does not change after it completes its simulation of the future state of $V^{\prime}$. This just means that I require that the simulating computer halts when it completes its simulation. I also require that the ID of the computer $N$ be replaced in toto (i.e., uniquely, with all other variables fixed to some predefined values) by the output of its simulation program. So for example, if $N$ is a multi-tape prefix TM, this means that when $N$ halts with the result of its simulation on its output tape, all the other tapes — the intermediate work tapes and the input tape — have been re-initialized to be all blanks.
@@
Note that Definition 1 does not require that the simulating computer $N$ calculates the future state of the universe $V^{\prime}$ by itself, independently of the initial state of the environment $W$ outside of $N$, i.e., independently of the value of $w_{0}$. (Formally, the LHS of Definition 1 is not independent of $\omega$, the initial state of $W$.) The reason for this flexibility is to allow the computer $N$ to retrieve the specific information it needs to perform its simulation of the dynamics of the specific state $v^{\prime}_{0}$ from those super-aliens who are running that computer $N$, and who exist in the environment of $N$, $W$. Note though that this freedom also allows the beings running the simulation computer $N$ to intervene on the dynamics of that computer, e.g., by overwriting the simulation program being run in the computer at any time they want. They can even “pull the plug early” on that computer, before it finishes its computation.
@@
Definition 1 allows both $w_{0}$ and $n_{0}$ to vary if we change $w^{\prime}_{0}$, even if $n^{\prime}_{0}$ is fixed. Often we are interested in a more restrictive notion of simulation, where for a fixed $n^{\prime}_{0}$, changing $w^{\prime}_{0}$ does not change $n_{0}$ (even though changing $w^{\prime}_{0}$ will of course change $w_{0}$ in general). Intuitively, this restrictive form of simulation corresponds to the case where the simulation program is fixed, reading in the precise initial state of the system it is simulating after it starts running. As an example, this form of simulation would be met if $N$ were a UTM, so that $n_{0}$ specifies the precise TM that $N$ is implementing, while $w^{\prime}_{0}$ is extra information that is subsequently “read in” by that TM which is specified in $n_{0}$.
We can formalize this restricted form of simulation with a simple extension of Definition 1.
@@
Suppose that $V$ simulates the evolution of $V^{\prime}$ for three functions ${\mathcal{T}},{\mathcal{W}},{\mathcal{N}}$. Then $V$ freely simulates the evolution of $V^{\prime}$ if for some specific fixed $n^{\prime}_{0}$, ${\mathcal{N}}(\Delta t^{\prime},w^{\prime}_{0},n^{\prime}_{0})$ is independent of $w^{\prime}_{0}$.
Finally, the definitions above only stipulate that the simulating computer $N$ eventually outputs the future state of $V^{\prime}$ at one specific time, $v_{\Delta t^{\prime}}$. It is straight-forward to extend these definitions to have $N$ output an entire trajectory of $L$ such future states instead. To do this we would replace the first arguments of ${\mathcal{T}},{\mathcal{W}}$ and ${\mathcal{N}}$ with a vector $\vec{\Delta t^{\prime}}\in{\mathbb{N}}^{L}$. We would also extend the definition of the evolution function, to have
@@
be the states of $V^{\prime}$ at the sequence of times $\vec{\Delta t^{\prime}}$ when it starts at time 00 with the state $w^{\prime}_{0},n^{\prime}_{0})$. Finally, we would modify the condition in Eq. 3 to say that $V$ trajectory-simulates $V^{\prime}$ if
@@ 2.4 The Physical Church-Turing Thesis
Even restricting consideration to computers that can be described using classical physics, there are many different semi-formal definitions of the PCT in the literature [50, 49, 23, 24, 3]. If we extend consideration to include quantum computers [46], there are even more definitions [6, 45].
Whether in fact our particular cosmological universe obeys the PCT, be it the classical or quantum PCT, has been subject to endless argument [2, 48, 6, 49, 47]. In particular, some researchers have designed purely theoretical, contrived physical systems that are uncomputable in some sense or other [51, 25, 54] (see also [19]). This work has resulted in attempts to define the PCT to exclude the case of physical systems whose future is uncomputable but which cannot be constructed by we humans in a finite amount of time. This amounts to tightening the PCT to concern not just what systems can be simulated, but rather what systems can be constructed and then simulated.
@@
(Note that this is the first fully general definition of the PCT, even applicable to universes whose laws of physics differ from ours.) By Definition 3, if the PCT holds for $V$, there must be a UTM that (halts and) outputs the vector value of $g(.,.,.)$, for all values of $g$’s arguments. (More precisely, there must be such a UTM that outputs the string $\langle g(.,.,.)\rangle$ for all values of $g$’s arguments if it receives the encoded version of those arguments as its input.)
All the analysis in this paper will assume that at least some of the universes being discussed obey the PCT. This means that the analysis below does not apply to any universes so many levels of computational power above our own that they can contain computers capable of super-Turing computation [3]. In particular, the analysis would not apply to any such universes that are simulating our universe.
Much of the earlier literature on the “physical Church-Turing thesis” accords a prominent role to humans, and their abilities (or lack thereof), e.g., in arbitrarily configuring the initial state of physical systems, or in observing their subsequent physical state. There is no such role in Definition 3. All the PCT means in this paper is that evolving the computational universe $V$ does not require a computational machine more powerful than a TM. In fact, it could be that that evolution can be calculated on a machine that is strictly weaker than a TM, e.g., a finite automaton. Furthermore, if $V$ is a sub-region of some cosmological universe, it could be that it requires machines more powerful than a TM to calculate the evolution of some other sub-region of that cosmological universe, different from $V$.
For these kinds of reasons, some readers might argue that Definition 3 doesn’t exactly capture any of the various properties that have been referred to as the “physical Church Turing thesis” in the literature. In some senses it more like one of the related concepts inspired by modern physics, e.g., some forms of ontic structural realism [28, 40, 39, 4] or the level IV multiverse [59, 60]. But for current purposes, we can ignore these semantic distinctions.
Finally, note that the set of spaces $W\times N$ is countably infinite, if we restrict attention to any and all finite $W$. Therefore the set of universes defined by the specification of such a space, together with an evolution function that obeys the PCT, is also countably infinite. Moreover, many of the considerations of the “simulation hypothesis” in the literature implicitly assume such a universe and evolution function. Finally, note that it is impossible to assign a uniform probability distribution to the set ${\mathbb{N}}$. This establishes the claim made in the introduction, that it is impossible to assign a uniform probability distribution to the kind of universes often considered in the literature on the simulation hypothesis.

### v3

@@ 2.1 Notation
My notation is conventional. The set of of all positive integers is $\mathbb{N}$}, and the set of non-negative integers is ${\mathbb{Z}}^{+}={\mathbb{N}}\cup\{0\}$. I write $|X|$ for the cardinality of any set $X$. In addition, for any set $X$ I write $X^{*}$ for the set of all finite strings of elements of $X$. Note that so long as $X$ is finite, $X^{*}$ is countable. As an important example, $\mathbb{B}$$^{*}$ is the set of all finite bit strings. Since that set can be bijectively mapped to $\mathbb{N}$}, I will follow convention and treat finite bit strings as positive integers and vice-versa, with the bijection implicit.
Appendix A reviews Turing machine theory, how each separate TM can be given its own index value from $\mathbb{N}$}, etc. As discussed there in detail, I write $T^{m}(x)$ for the (possibly partial) function given by running the TM with index $m$ on input $x$ until it halts (with the function undefined for input $x$ if it does not halt for that input). So $m$ ranges over all natural numbers $\mathbb{N}$}. Often I assume, implicitly or otherwise, that certain quantities are in the form of prefix-encoded bit strings. In particular, I use some standard bijective encoding function of all tuples of bit strings into a single bit string, indicated using angle brackets, $.,.\langle$$\rangle$$\langle$$\rangle$, etc. I assume that this encoding is non-decreasing in the number of arguments, i.e., for any finite set of bit strings $\{b(1),b(2),\ldots,b(m)\}$, the length of the bit string $\langle b(1),b(2),\ldots,b(m-1)\rangle$ is not greater than that of $\langle b(1),b(2),\ldots,b(m)\rangle$.
Other important definitions and notations are in Appendix A , e.g., of partial functions, and of computable functions (here always assumed to be total). I also review the definition of “computational universality” in that appendix. In addition, there I review the definitions of a universal TM (UTM), and a prefix-free TM (the implementation of TMs I will often be assume in this paper, implicitly or otherwise). I also define the instantaneous description (ID) of a Turing machine (TM) there. In the main body of this paper I will assume that the non-blank alphabet of the TM, $\Lambda$, is just $\mathbb{B}$}. So as described in Appendix A, we can take the state of the tape to be a finite string in $\mathbb{B}$$^{*}$, even though strictly speaking the definition of TMs assumes infinitely long tapes.
@@ 2.2 Framework for analyzing CS theory of the simulation hypothesis
I will also mostly be interested in cases where the computer simulates the evolution of the state of the universe external to the computer (the environment of the computer) and/or its own evolution. To simplify the analysis of such scenarios, I will assume the state space of the environment is finite (though arbitrarily large), to ensure that its state at a particular time can be appropriately encoded on the input of a UTM. (This assumption can be relaxed, at the cost of complicating the analysis.) This leads me to write the state space of a (computational) universe as
@@
where $W$ is initially finite and cannot be extended dynamically, whereas $N$ is countably infinite.
$W$ is the set of possible states of the computational universe external to a computer that it contains. $N$ is the set of possible states of that computer. I will parameterize $N$ as $\underline{N}\times R$. $R$ is a finite set that represents the internal variables of the computer. So for example, in a conventionally represented TM, $R$ would be the state of the TM’s head, the position(s) of its pointer(s) on the tape(s), etc. $\underline{N}={\mathbb{B}}^{*}$ is instead the state of the tape(s) in the case that the computer is represented as a TM, or it is the state of the memory in the case that the computer is represented as a RAM machine, or is represented as an appropriately modified laptop computer.  [^9] The reason for writing the countably infinite space $\underline{N}$ as the set of finite bit strings is so that we can physically implement it as a physical system that initially has a finite state space, but whose state space can be expanded by an arbitrarily large finite amount an arbitrary number of times as the system evolves. See Section 2.6., [^10] Strictly speaking, in a TM the pointer variables can have any value in $\mathbb{N}$}, since the TM’s head can point to an arbitrary position on the (countably infinite) tape. This would seem to contradict the definition that the pointer’s value lies in the finite space $R$. This technical concern can be addressed several ways. Perhaps most simply, we could just have $R$ infinitely extendable. From now on to simplify the language, I will refer to the “tape” of the UTM, singular, even if I am representing the case of a multi-tape TM. I will write the elements of $\underline{N}$ as $\underline{n}$.
@@
To simplify the analysis in this paper, I do not assume that the models of physical universes that I investigate capture those universes in toto, e.g., if the state spaces of those entire universes are actually uncountably infinite. Similarly, I will not assume that a simulation running in a computer inside a universe models the dynamics of an entire universe. Rather throughout this paper the expression “universe” should be understand as shorthand for “possibly coarse-grained subset of a universe”. However, I do assume that the universe obeys deterministic dynamics, be it over the original state space (assuming it is countable) or some coarse-grained version of that state space. As an example, in Section 2.6 I present a detailed example of what such a “coarse-grained subset of a universe” could be for our particular physical universe.
I write the evolution of the universe from an initial, time-0 state $(w_{0},n_{0})$ to a time $t$ state as a vector-valued evolution function $g$ of the initial state of that universe,
@@
Note that the image of $g$ is the Cartesian product $W\times N=W\times(\underline{N}\times R)$. I use the usual notation for components of vector-valued functions. So in particular, $g(t,w_{0},n_{0})_{\underline{N}}$ is the $\underline{N}$-component of $g(t,w_{0},n_{0})$.
@@
As shorthand, below I will often “propagate” accents on symbols concerning computational universes. For example, when I say below that “$V$ simulates $V^{\prime}$”, I will implicitly assume $V=W\times N$, and that $V^{\prime}=W^{\prime}\times N^{\prime}$”. All universes, with whatever accents, are assumed to be defined for the set of times ${\mathbb{Z}}^{+}=\{0,1,\ldots\}$.
@@ 2.3 What it means for one universe to simulate another
The term “simulation” has not been given a formal definition in any of the previous literature on the simulation hypothesis. One of the effects of this lack of formal rigor has been to prevent the exploitation of CS theory to investigate the simulation hypothesis. Another effect is to obscure some important distinctions between different types of simulation. (These are introduced below.) More directly though, this lack of formal rigor encourages confusion when one tries to exploit CS theory to understand the simulation hypothesis. For example, “simulation” (and the associated term “bisimulation”) already has a formal definition in the CS theory of state transition systems [68]. However, this definition from CS theory does not describe what “simulation” is loosely understood to mean in the context of the simulation hypothesis. In this subsection I fill this hole in the literature, providing a fully formal definition of “simulation” as the term is used in the simulation hypothesis.
Intuitively, the idea is that $V$ “simulates” $V^{\prime}$ if for all initial states $v^{\prime}_{0}$ of $V$, and all amounts of time $\Delta t^{\prime}$ into the future of $V^{\prime}$ that we might want to simulate the dynamics of $V^{\prime}$, there is some associated initial state of $N$ and of $W$, together with a time into the future of the simulating universe, $\Delta t$, such that the state of the computer $N$ at that future time $\Delta t$ gives us the desired state of $V^{\prime}$ at its future time $\Delta t^{\prime}$. That triple of associated initial state of $N$ and of $W$ and a time $\Delta t$ are functions of $v^{\prime}_{0},\Delta t^{\prime}$.
Writing $v^{\prime}_{0}$ as $(w^{\prime}_{0},n^{\prime}_{0})$, this requirement can be formalized as follows:
@@
Note that the LHS of Eq. 3 is the second of the two components of the vector-valued function $g$, while the RHS is an encoding of both components of $g^{\prime}$ into a single variable. Note also the requirement in Definition 1 that Eq. 3 hold for all $t\geq\tau$, and so the state of the simulating computer $N$ does not change after it completes its simulation of the future state of $V^{\prime}$. This just means that I require that the simulating computer halts when it completes its simulation.
To simplify the analysis, I also require that the ID of the computer $N$ be replaced in toto (i.e., uniquely, with all other variables fixed to some predefined, post-halting values) by the output of its simulation program. So for example, if $N$ is a multi-tape prefix TM, this means that when $N$ halts with the result of its simulation on its output tape, all the other tapes — the intermediate work tapes and the input tape — have been re-initialized to be all blanks.
@@
Note that Definition 1 does not require that the simulating computer $N$ calculates the future state of the universe $V^{\prime}$ by itself, independently of the initial state of the environment $W$ outside of $N$, i.e., independently of the value of $w_{0}$. (Formally, the function $g$ on the LHS of Eq. 3 is not independent of $\omega$, the initial state of $W$.) The reason for this flexibility is to allow the computer $N$ to retrieve the specific information it needs to perform its simulation of the dynamics of the specific state $v^{\prime}_{0}$ from those super-aliens who are running that computer $N$, and who exist in the environment of $N$, $W$.
In addition, since there are no restrictions in Definition 1 on the functions ${\mathcal{T}},{\mathcal{W}}$ and $\mathcal{N}$}, the second and third arguments of $g$ in Eq. 3 can vary arbitrarily. This means there are no a priori restrictions on how the full universe $V$ can evolve. As a result, Eq. 3 allows any putative beings running the simulation computer $N$ to intervene on the dynamics of that computer after it has started running, e.g., by overwriting the simulation program being run, in a completely arbitrary fashion. They can even “pull the plug early” on that computer, before it finishes its computation.
@@
Definition 1 allows both $w_{0}$ and $n_{0}$ to vary if we change $w^{\prime}_{0}$, even if $n^{\prime}_{0}$ is fixed. Often we are interested in a more restrictive notion of simulation, where for a fixed $n^{\prime}_{0}$, changing $w^{\prime}_{0}$ does not change $n_{0}$, even though changing $w^{\prime}_{0}$ is allowed to change $w_{0}$ in general. Intuitively, this restrictive form of simulation corresponds to the case where the simulation program $n_{0}$ is fixed by $n^{\prime}_{0}$, and that simulating computer “reads in” the precise initial state of the entire universe it is simulating at some iteration $t>0$ after it starts running, by appropriately interacting with the state of its own external universe at that time, $w_{t}$. As an example, this form of simulation would be met if $N$ were a UTM, so that $n_{0}$ specifies the precise TM that $N$ is implementing, while $w^{\prime}_{0}$ is extra information that is “read in” at a later time $t$ by that TM $n_{0}$, from the associated state $w_{t}$.
We can formalize this special case of simulation with a simple extension of Definition 1:
@@
Suppose that $V$ simulates the evolution of $V^{\prime}$ for three functions ${\mathcal{T}},{\mathcal{W}},{\mathcal{N}}$. Then $V$ freely simulates $V^{\prime}$ for $n^{\prime}_{0},\Delta t^{\prime}$, if ${\mathcal{N}}(\Delta t^{\prime},w^{\prime}_{0},n^{\prime}_{0})$ is independent of $w^{\prime}_{0}$.
Note that this definition allows the evolution of $n^{\prime}$ to depend on $w^{\prime}_{0}$, i.e., it does not restrict the interactions between the computer being simulated and its external world. Rather the restriction only affects initialization details concerning the universe doing the simulating.
Finally, the definitions above only stipulate that the simulating computer $N$ eventually outputs the future state of $V^{\prime}$ at one specific time, $v_{\Delta t^{\prime}}$. While sufficient for this paper, it is worth pointing out that we could extend these definitions to instead have $N$ output an entire trajectory of $L$ such future states for some finite number $L$. To do this we would replace the first arguments of ${\mathcal{T}},{\mathcal{W}}$ and $\mathcal{N}$} with a vector $\vec{\Delta t^{\prime}}\in{\mathbb{N}}^{L}$. We would also extend the definition of the evolution function, to have
@@
be the states of $V^{\prime}$ at an entire sequence of times $\vec{\Delta t^{\prime}}$ which arise if it starts at time 0 with the state $(w^{\prime}_{0},n^{\prime}_{0})$. Finally, we would modify the condition in Eq. 3 to say that $V$ trajectory-simulates $V^{\prime}$ if
@@ 2.4 The Physical Church-Turing Thesis
Even restricting consideration to computers that can be described using classical physics, there are many different semi-formal definitions of the PCT in the physics literature [55, 54, 24, 25, 2]. If we extend consideration to include quantum computers [51], there are even more definitions [6, 50].
Whether in fact our particular cosmological universe obeys the PCT, be it the classical or quantum PCT, has been subject to endless argument [1, 53, 6, 54, 52]. In particular, some researchers have designed purely theoretical, contrived physical systems that are uncomputable in some sense or other [57, 26, 59] (see also [20]). This work has resulted in attempts to define the PCT to exclude the case of physical systems whose future is uncomputable but which cannot be constructed by us humans in a finite amount of time. This amounts to tightening the PCT to concern not just what systems can be simulated, but rather what systems can be constructed and then simulated.
@@
Definition 3 is the first fully general definition of the PCT, even applicable to universes whose laws of physics differ from ours.
By Definition 3, if the PCT holds for $V$, there must be a UTM that (halts and) outputs the vector value of $g(.,.,.)$, for all values of $g$’s arguments. More precisely, there must be such a UTM that outputs the string $\langle g(.,.,.)\rangle$ for all values of $g$’s arguments if it receives the encoded version of those arguments as its input.
The analysis below implicitly assumes that all of the universes being discussed obey the PCT. This means that the analysis below does not apply to any universes that can contain computers capable of super-Turing computation [2]. In particular, the analysis would not apply to any such universes that are simulating our universe.
Much of the earlier literature on the “physical Church-Turing thesis” accords a prominent role to humans, and their abilities (or lack thereof), e.g., in arbitrarily configuring the initial state of physical systems, or in observing their subsequent physical state. There is no such role in Definition 3. All the PCT means in this paper is that evolving the computational universe $V$ does not require a computational machine more powerful than a TM. In fact, it could be that that evolution can be calculated on a machine that is strictly weaker than a TM, e.g., a finite automaton, and it would still obey Definition 3. Furthermore, Definition 3 could still apply if $V$ is a only sub-region of some cosmological universe, and machines more powerful than a TM are required to calculate the evolution of some other sub-region of that cosmological universe, different from $V$.
For these kinds of reasons, some readers might argue that Definition 3 doesn’t exactly capture any of the various properties that have been referred to as the “physical Church Turing thesis” in the literature. In some senses it more like one of the related concepts inspired by modern physics, e.g., some forms of ontic structural realism [29, 44, 43, 4] or the level IV multiverse [63, 64]. But for current purposes, we can ignore these terminological distinctions.
Finally, note that the set of spaces $W\times N$ is countably infinite, if we restrict attention to any and all finite $W$. Therefore the set of universes defined by the specification of such a space, together with an evolution function that obeys the PCT, is also countably infinite. Moreover, many of the considerations of the “simulation hypothesis” in the literature implicitly assume such a universe and evolution function. However, it is impossible to assign a uniform probability distribution to the set $\mathbb{N}$}. This establishes the claim made in Appendix E, that it is impossible to assign a uniform probability distribution to the kind of universes often considered in the literature on the simulation hypothesis. This contradicts an assumption made in much of the philosophy literature on the simulation hypothesis.